# On subgroups with narrow Schreier graphs

### v2

@@ 3 Narrow spaces
Since $H$ has at least two relative ends and bounded packing, there is an action of $G$ on an unbounded finite dimensional CAT(0) cube complex $C$ with a single orbit of hyperplanes $G\cdot\operatorname{\mathfrak{h}}$ such that $H$ has finite index in the $G$-stabiliser of $\operatorname{\mathfrak{h}}$, denoted $\operatorname{Stab}_{G}(\operatorname{\mathfrak{h}})$ (see Section 2). By Proposition 3.4 and since being a virtual fiber subgroup is closed under commensurability, we can assume without loss of generality that $H=\operatorname{Stab}_{G}(\operatorname{\mathfrak{h}})$. By Proposition 2.1, the action of $G$ on $C$ is essential and there is a hyperbolic element $\gamma\in G$ which skewers $\mathfrak{h}$. Label the halfspaces of $\operatorname{\mathfrak{h}}$ and replace $\gamma$ with a proper power if necessary so that $\gamma\operatorname{\mathfrak{h}}^{+}\subsetneq\operatorname{\mathfrak{h}}^{+}$. No power of $\gamma$ is contained in $H$ so $\rho_{H}(\operatorname{\langle}\gamma\operatorname{\rangle})$ is infinite. If $\mu$ is the word length of $\gamma$ with respect to $S\cup S^{-1}$ then $\rho_{H}(g\operatorname{\langle}\gamma\operatorname{\rangle})$ is $\mu$-coarsely connected for all $g\in G$. Therefore there exists $N\in\mathbb{N}$ and $k_{1},\dots,k_{N}\in G$ such that the $\rho_{H}$-images of $\{k_{i}\operatorname{\langle}\gamma\operatorname{\rangle}\}$ are infinite and pairwise distinct and, for all $k\in G$, if $\rho_{H}(k\operatorname{\langle}\gamma\operatorname{\rangle})$ is infinite then $\rho_{H}(k\operatorname{\langle}\gamma\operatorname{\rangle})=\rho_{H}(k_{i}\operatorname{\langle}\gamma\operatorname{\rangle})$ for some $i$.
@@
Let $k\in G$ be such that $k\operatorname{\mathfrak{h}}\cap\operatorname{\mathfrak{h}}=\emptyset$. Claim 1 implies that $\rho_{H}(k\operatorname{\langle}\gamma\operatorname{\rangle})$ is infinite, so there is $i\in\{1,\dots,N\}$ such that $Hk\operatorname{\langle}\gamma\operatorname{\rangle}=Hk_{i}\operatorname{\langle}\gamma\operatorname{\rangle}$. In particular, there exists $n\in\mathbb{Z}$ and $h\in H$ such that $k=hk_{i}\gamma^{n}$. Thus if $\ell$ is an axis for $\gamma$ then $d(\operatorname{\mathfrak{h}},k\ell)=d(\operatorname{\mathfrak{h}},hk_{i}\gamma^{n}\ell)=d(\operatorname{\mathfrak{h}},k_{i}\ell)$. Let $M\coloneqq\max\{d(\operatorname{\mathfrak{h}},k_{i}\ell):1\leq i\leq N\}$.
@@
Let $g,g^{\prime}\in G$ be such that $\{\operatorname{\mathfrak{h}},g\operatorname{\mathfrak{h}},g^{\prime}\operatorname{\mathfrak{h}}\}$ is a facing triple. Then $\gamma$ cannot skewer both $g\operatorname{\mathfrak{h}}$ and $g^{\prime}\operatorname{\mathfrak{h}}$ so, up to relabelling, we can assume that $\ell\cap g\operatorname{\mathfrak{h}}=\emptyset$. This implies that $\operatorname{\mathfrak{h}}$ and $\ell$ are both contained in a halfspace bounded by $g\operatorname{\mathfrak{h}}$, which we denote by $(g\operatorname{\mathfrak{h}})^{+}$. Let $\psi=g\gamma g^{-1}$ and note that $\psi$ skewers $g\operatorname{\mathfrak{h}}$. Let $n\in\mathbb{Z}$ be such that $\psi^{n}(g\operatorname{\mathfrak{h}})^{-}\subsetneq(g\operatorname{\mathfrak{h}})^{-}$. Then for all $i\in\mathbb{N}$,
@@ 5 Reducing a Euclidean cube complex to a line
Let $\operatorname{\mathfrak{k}}$ be a hyperplane which intersects $\operatorname{\mathfrak{h}}$ but not $\gamma\operatorname{\mathfrak{h}}$. If $\operatorname{\mathfrak{k}}$ separates $\gamma\operatorname{\mathfrak{h}}$ from $\gamma^{-n}\operatorname{\mathfrak{h}}$ for some $n\in\mathbb{N}$ then $\gamma$ skewers $\operatorname{\mathfrak{k}}$. So suppose that $\operatorname{\mathfrak{k}}$ does not separate $\gamma\operatorname{\mathfrak{h}}$ from any $\gamma^{-n}\operatorname{\mathfrak{h}}$. Since there are no facing triples, $\operatorname{\mathfrak{k}}$ must intersect $\gamma^{-n}\operatorname{\mathfrak{h}}$ for all $n\in\mathbb{N}$. Therefore for all $n\geq 1$, $\gamma^{n}\operatorname{\mathfrak{k}}$ intersects both $\operatorname{\mathfrak{h}}$ and $\gamma\operatorname{\mathfrak{h}}$ while $\gamma^{-n}\operatorname{\mathfrak{k}}$ intersects neither $\operatorname{\mathfrak{h}}$ nor $\gamma\operatorname{\mathfrak{h}}$. Similarly, for any $\operatorname{\mathfrak{k}}\in\mathcal{H}(\operatorname{\mathfrak{h}})\Delta\mathcal{H}(\gamma\operatorname{\mathfrak{h}})$ and $n\neq 0$, $\gamma^{n}\operatorname{\mathfrak{k}}\notin\mathcal{H}(\operatorname{\mathfrak{h}})\Delta\mathcal{H}(\gamma\operatorname{\mathfrak{h}})$. This implies in particular that any $\operatorname{\mathfrak{k}}\in\mathcal{H}(\operatorname{\mathfrak{h}})\Delta\mathcal{H}(\gamma\operatorname{\mathfrak{h}})$ is not stabilised by any power of $\gamma$ and, by Lemma 2.3, must therefore be skewered by $\gamma$.
@@
Let $\mathcal{H}$ be a (necessarily finite) set of pairwise intersecting hyperplanes of $Y$ and $g\in\Gamma$ be an isometry which skewers each $\operatorname{\mathfrak{h}}\in\mathcal{H}$. Label the halfspaces and replace $g$ with a proper power if necessary so that $g\operatorname{\mathfrak{h}}^{+}\subsetneq\operatorname{\mathfrak{h}}^{+}$ for each $\operatorname{\mathfrak{h}}\in\mathcal{H}$. Suppose that there exists a hyperplane $\operatorname{\mathfrak{k}}$ of $Y$ which is not skewered by $g$ and let $k\in\Gamma$ be such that $k\operatorname{\mathfrak{k}}^{+}\subsetneq\operatorname{\mathfrak{k}}^{+}$. By Lemma 2.3, there exists $n_{1}\in\mathbb{N}$ such that $g^{n_{1}}\operatorname{\mathfrak{k}}^{+}=\operatorname{\mathfrak{k}}^{+}$. Moreover, there exists $m\in\mathbb{N}$ such that, for each $\operatorname{\mathfrak{h}}\in\mathcal{H}$, either $k^{m}\operatorname{\mathfrak{h}}^{+}=\operatorname{\mathfrak{h}}^{+}$, $k^{m}\operatorname{\mathfrak{h}}^{+}\subsetneq\operatorname{\mathfrak{h}}^{+}$ or $k^{m}\operatorname{\mathfrak{h}}^{-}\subsetneq\operatorname{\mathfrak{h}}^{-}$. In the first two cases, $g^{n_{1}n}k^{m}\operatorname{\mathfrak{h}}^{+}\subseteq g^{n_{1}n}\operatorname{\mathfrak{h}}^{+}\subsetneq\operatorname{\mathfrak{h}}^{+}$ for all $n\in\mathbb{N}$. In the latter case, note that $d_{Haus}(\operatorname{\mathfrak{h}},k^{m}\operatorname{\mathfrak{h}})<\infty$ by Lemma 5.1 and, for any $n\in\mathbb{N}$ such that $d(g^{nn_{1}}\operatorname{\mathfrak{h}},\operatorname{\mathfrak{h}})>d_{Haus}(\operatorname{\mathfrak{h}},k^{m}\operatorname{\mathfrak{h}})$, we have that $g^{n_{1}n}k^{m}\operatorname{\mathfrak{h}}^{+}\subsetneq\operatorname{\mathfrak{h}}^{+}$. Since there are finitely many hyperplanes in $\mathcal{H}$, we can choose $n_{2}\in\mathbb{N}$ sufficiently large so that $g^{n_{1}n_{2}}k^{m}\operatorname{\mathfrak{h}}^{+}\subsetneq\operatorname{\mathfrak{h}}^{+}$ for all $\operatorname{\mathfrak{h}}\in\mathcal{H}$. We also have that $g^{n_{1}n_{2}}k^{m}\operatorname{\mathfrak{k}}^{+}\subsetneq g^{n_{1}n_{2}}\operatorname{\mathfrak{k}}^{+}=\operatorname{\mathfrak{k}}^{+}$ so the element $g^{\prime}\coloneqq g^{n_{1}n_{2}}k^{m}$ skewers every hyperplane in $\mathcal{H}^{\prime}\coloneqq\mathcal{H}\cup\{\operatorname{\mathfrak{k}}\}$. Moreover, for each $\operatorname{\mathfrak{h}}\in\mathcal{H}$, $\operatorname{\mathfrak{k}}$ cannot separate hyperplanes in $\operatorname{\langle}g\operatorname{\rangle}\cdot\operatorname{\mathfrak{h}}$ without being skewered by $g$. By the lack of facing triples, this implies that $\operatorname{\mathfrak{k}}$ must intersect $g^{n}\operatorname{\mathfrak{h}}$ for some $n$. It then follows from Lemma 5.1 that $\operatorname{\mathfrak{k}}$ intersects $g^{n}\operatorname{\mathfrak{h}}$ for all $n$. In particular, $\operatorname{\mathfrak{k}}$ intersects every hyperplane in $\mathcal{H}$, so $\mathcal{H}^{\prime}$ is pairwise intersecting and $|\mathcal{H}^{\prime}|\leq\dim(Y)$. If there exists a hyperplane $\operatorname{\mathfrak{k}}^{\prime}$ which is not skewered by $g^{\prime}$, then the above argument applied to $\mathcal{H}^{\prime},g^{\prime}$ and $\operatorname{\mathfrak{k}}^{\prime}$ implies that ′′$\mathcal{H}^{\prime\prime}\coloneqq\mathcal{H}^{\prime}\cup\{\operatorname{\mathfrak{k}}^{\prime}\}$ is pairwise intersecting and there exists $g^{\prime\prime}\in\Gamma$ which skewers every hyperplane in $\mathcal{H}^{\prime\prime}$. We can only iterate this process at most $\dim(Y)-1$ times until we obtain the desired element $\gamma$. ∎
@@
is well defined and, by [Cor16, Proposition 4.H.1], it is a homomorphism.
Clearly $K\leq\ker(tr_{M^{+}})$. Let us show that $K$ has finite index in $\ker(tr_{M^{+}})$. Given a hyperplane $\operatorname{\mathfrak{k}}$ of $Y$, define $L_{\operatorname{\mathfrak{k}}}$ as follows. If $\operatorname{\mathfrak{k}}$ does not separate $\gamma^{n_{1}}\operatorname{\mathfrak{h}}$ from $\gamma^{n_{2}}\operatorname{\mathfrak{h}}$ for any $n_{1},n_{2}\in\mathbb{Z}$ then $L_{\operatorname{\mathfrak{k}}}\coloneqq 0$. Otherwise
@@
It remains to show that $\Gamma^{\prime}$ has finite index in $\Gamma$. Let $R\coloneqq d_{Haus}(\operatorname{\mathfrak{h}},\gamma\operatorname{\mathfrak{h}})<\infty$, let $\ell$ be a combinatorial geodesic axis for $\gamma$ and let $\sigma\subseteq\ell$ be a segment of length $R$. Let $1\leq R^{\prime}\leq R$ be such that exactly $R^{\prime}$ of the edges of $\sigma$ are dual to hyperplanes in the $\Gamma$-orbit of $\operatorname{\mathfrak{h}}$ and let $g_{1},\dots,g_{R^{\prime}}\in\Gamma$ be such that each hyperplane in $\Gamma\cdot\operatorname{\mathfrak{h}}$ which intersects $\sigma$ is of the form $g_{i}\operatorname{\mathfrak{h}}$ for some $i$. Then, for any $g\in\Gamma$, there is some $n\in\mathbb{Z}$ and $i\in\{1,\dots,R\}$ such that $g\gamma^{n}\operatorname{\mathfrak{h}}=g_{i}\operatorname{\mathfrak{h}}$. Therefore $g_{i}^{-1}g\gamma^{n}\in K$ and $g_{i}^{-1}g\in K\gamma^{-n}\subseteq\Gamma^{\prime}$. Hence $\Gamma^{\prime}$ has index at most $R$ in $\Gamma$. ∎
@@
If we assume, as in the proof of Proposition 4.2, that $\Gamma$ and $\operatorname{Stab}_{\Gamma}(\operatorname{\mathfrak{h}})$ are finitely generated and $\operatorname{Stab}_{\Gamma}(\operatorname{\mathfrak{h}})\operatorname{\setminus}\Gamma$ is narrow, then there is a much shorter proof:
For each $n\in\mathbb{N}$ let $\mathcal{H}_{n}$ denote the set of hyperplanes whose distance from $\operatorname{\mathfrak{h}}$ is exactly $n$. Then the hyperplanes of $\mathcal{H}_{n}$ contained in $\operatorname{\mathfrak{h}}^{-}$ are pairwise intersecting, as are those contained in $\operatorname{\mathfrak{h}}^{+}$. Therefore $|\mathcal{H}_{n}|\leq 2\dim(Y)$ and, since $\operatorname{Stab}_{\Gamma}(\operatorname{\mathfrak{h}})$ acts on $\mathcal{H}_{n}$, there is a homomorphism $f_{n}:\operatorname{Stab}_{\Gamma}(\operatorname{\mathfrak{h}})\rightarrow S(2\dim(C))$, where $S(2\dim(C))$ is the symmetric group on $2\dim(C)$ elements and $\ker f_{n}$ acts trivially on $\mathcal{H}_{n}$. The kernel $\ker f_{n}$ has index at most $(2\dim(C))!$ in $\operatorname{Stab}_{\Gamma}(\operatorname{\mathfrak{h}})$ and, since $\operatorname{Stab}_{\Gamma}(\operatorname{\mathfrak{h}})$ is finitely generated, there are finitely many such subgroups. Thus the intersection $K\coloneqq\cap_{n\in\mathbb{N}}\ker f_{n}$ is a finite intersection of finite index normal subgroups of $\operatorname{Stab}_{\Gamma}(\operatorname{\mathfrak{h}})$ and is itself a normal subgroup of $\operatorname{Stab}_{\Gamma}(\operatorname{\mathfrak{h}})$ with finite index. Let $g\in\Gamma$ be such that $g\operatorname{\mathfrak{h}}^{+}\subsetneq\operatorname{\mathfrak{h}}^{+}$. Then $K$ fixes $g^{n}\operatorname{\mathfrak{h}}$ for all $n\in\mathbb{Z}$, so $K$ has infinite index in its normaliser. Since $\operatorname{Stab}_{\Gamma}(\operatorname{\mathfrak{h}})\leq N(K)$, this implies that $[\Gamma:N(K)]<\infty$ by Proposition 3.5. The quotient group $N(K)/K$, equipped with any proper word metric, is quasi-isometric to $K\operatorname{\setminus}\Gamma$, which is a finite index cover of $\operatorname{Stab}_{\Gamma}(\operatorname{\mathfrak{h}})\operatorname{\setminus}\Gamma$. Since narrow graphs have finitely many ends, it follows that $N(K)/K$ has $2\leq e<\infty$ ends but, since $N(K)/K$ is a group, $e$ must equal $2$ and $N(K)/K$ must be virtually cyclic [Hop44].

### v3

@@ 3 Narrow spaces
Since $H$ has at least two relative ends and bounded packing, there is an action of $G$ on an unbounded finite dimensional CAT(0) cube complex $C$ with a single orbit of hyperplanes $G\cdot\operatorname{\mathfrak{h}}$ such that $H$ has finite index in the $G$-stabiliser of $\operatorname{\mathfrak{h}}$, denoted $\operatorname{Stab}_{G}(\operatorname{\mathfrak{h}})$ (see Section 2). By Proposition 3.5 and since being a virtual fiber subgroup is closed under commensurability, we can assume without loss of generality that $H=\operatorname{Stab}_{G}(\operatorname{\mathfrak{h}})$. By Proposition 2.1, the action of $G$ on $C$ is essential and there is a hyperbolic element $\gamma\in G$ which skewers $\mathfrak{h}$. Label the halfspaces of $\operatorname{\mathfrak{h}}$ and replace $\gamma$ with a proper power if necessary so that $\gamma\operatorname{\mathfrak{h}}^{+}\subsetneq\operatorname{\mathfrak{h}}^{+}$. No non-zero power of $\gamma$ is contained in $H$ so $\rho_{H}(\operatorname{\langle}\gamma\operatorname{\rangle})$ is infinite. If $\mu$ is the word length of $\gamma$ with respect to $S\cup S^{-1}$ then $\rho_{H}(g\operatorname{\langle}\gamma\operatorname{\rangle})$ is $\mu$-coarsely connected for all $g\in G$. Therefore there exists $N\in\mathbb{N}$ and $k_{1},\dots,k_{N}\in G$ such that the $\rho_{H}$-images of $\{k_{i}\operatorname{\langle}\gamma\operatorname{\rangle}\}$ are infinite and pairwise distinct and, for all $k\in G$, if $\rho_{H}(k\operatorname{\langle}\gamma\operatorname{\rangle})$ is infinite then $\rho_{H}(k\operatorname{\langle}\gamma\operatorname{\rangle})=\rho_{H}(k_{i}\operatorname{\langle}\gamma\operatorname{\rangle})$ for some $i$.
@@
Let $k\in G$ be such that $k\operatorname{\mathfrak{h}}\cap\operatorname{\mathfrak{h}}=\emptyset$. Claim 1 implies that $\rho_{H}(k\operatorname{\langle}\gamma\operatorname{\rangle})$ is infinite, so there is $i\in\{1,\dots,N\}$ such that $Hk\operatorname{\langle}\gamma\operatorname{\rangle}=Hk_{i}\operatorname{\langle}\gamma\operatorname{\rangle}$. In particular, there exists $n\in\mathbb{Z}$ and $h\in H$ such that $k=hk_{i}\gamma^{n}$. Thus if $\ell$ is an $\ell^{2}$-geodesic axis for $\gamma$ then $d(\operatorname{\mathfrak{h}},k\ell)=d(\operatorname{\mathfrak{h}},hk_{i}\gamma^{n}\ell)=d(\operatorname{\mathfrak{h}},k_{i}\ell)$. Let $M\coloneqq\max\{d(\operatorname{\mathfrak{h}},k_{i}\ell):1\leq i\leq N\}$.
@@
Suppose that there exist $g,g^{\prime}\in G$ such that $\{\operatorname{\mathfrak{h}},g\operatorname{\mathfrak{h}},g^{\prime}\operatorname{\mathfrak{h}}\}$ is a facing triple. By [CS11, Lemma 2.3], $\ell$ intersects $\operatorname{\mathfrak{h}}$ at a single point and the complementary components of $\ell$ lie in different halfspaces of $\operatorname{\mathfrak{h}}$. This implies that $\ell$ is not contained in either $g\operatorname{\mathfrak{h}}$ or $g^{\prime}\operatorname{\mathfrak{h}}$. Therefore if $\ell$ intersects $g\operatorname{\mathfrak{h}}$ (resp. $g^{\prime}\operatorname{\mathfrak{h}}$) then the complementary components of $\ell$ lie in different halfspaces of $g\operatorname{\mathfrak{h}}$ (resp. $g^{\prime}\operatorname{\mathfrak{h}}$). Thus if $\ell$ intersects both $g\operatorname{\mathfrak{h}}$ and $g^{\prime}\operatorname{\mathfrak{h}}$ then $\{\operatorname{\mathfrak{h}},g\operatorname{\mathfrak{h}},g^{\prime}\operatorname{\mathfrak{h}}\}$ cannot be a facing triple. We can therefore assume, up to relabelling $g$ and $g^{\prime}$ if necessary, that $\ell\cap g\operatorname{\mathfrak{h}}=\emptyset$. This implies that $\operatorname{\mathfrak{h}}$ and $\ell$ are both contained in a halfspace bounded by $g\operatorname{\mathfrak{h}}$, which we denote by $(g\operatorname{\mathfrak{h}})^{+}$. Let $\psi=g\gamma g^{-1}$ and note that $\psi$ skewers $g\operatorname{\mathfrak{h}}$. Let $n\in\mathbb{Z}$ be such that $\psi^{n}(g\operatorname{\mathfrak{h}})^{-}\subsetneq(g\operatorname{\mathfrak{h}})^{-}$. Then for all $i\in\mathbb{N}$,
@@ 5 Reducing a Euclidean cube complex to a line
Let $\operatorname{\mathfrak{k}}$ be a hyperplane which intersects $\operatorname{\mathfrak{h}}$ but not $\gamma\operatorname{\mathfrak{h}}$. If $\operatorname{\mathfrak{k}}$ separates $\gamma\operatorname{\mathfrak{h}}$ from $\gamma^{-n}\operatorname{\mathfrak{h}}$ for some $n\in\mathbb{N}$ then $\gamma$ skewers $\operatorname{\mathfrak{k}}$. So suppose that $\operatorname{\mathfrak{k}}$ does not separate $\gamma\operatorname{\mathfrak{h}}$ from any $\gamma^{-n}\operatorname{\mathfrak{h}}$. Since there are no facing triples, $\operatorname{\mathfrak{k}}$ must intersect $\gamma^{-n}\operatorname{\mathfrak{h}}$ for all $n\in\mathbb{N}$. Therefore for all $n\geq 1$, $\gamma^{n}\operatorname{\mathfrak{k}}$ intersects both $\operatorname{\mathfrak{h}}$ and $\gamma\operatorname{\mathfrak{h}}$ while $\gamma^{-n}\operatorname{\mathfrak{k}}$ intersects neither $\operatorname{\mathfrak{h}}$ nor $\gamma\operatorname{\mathfrak{h}}$. Similarly, for any $\operatorname{\mathfrak{k}}\in\mathcal{H}(\operatorname{\mathfrak{h}})\Delta\mathcal{H}(\gamma\operatorname{\mathfrak{h}})$, either $\gamma$ skewers $\operatorname{\mathfrak{k}}$ or for all $n\neq 0$ we have $\gamma^{n}\operatorname{\mathfrak{k}}\notin\mathcal{H}(\operatorname{\mathfrak{h}})\Delta\mathcal{H}(\gamma\operatorname{\mathfrak{h}})$. This implies in particular that any $\operatorname{\mathfrak{k}}\in\mathcal{H}(\operatorname{\mathfrak{h}})\Delta\mathcal{H}(\gamma\operatorname{\mathfrak{h}})$ is not stabilised by any power of $\gamma$ and, by Lemma 2.3, must therefore be skewered by $\gamma$.
@@
Let $\mathcal{H}$ be a (necessarily finite) set of pairwise intersecting hyperplanes of $Y$ and $g\in\Gamma$ be an isometry which skewers each $\operatorname{\mathfrak{h}}\in\mathcal{H}$. Label the halfspaces and replace $g$ with a proper power if necessary so that $g\operatorname{\mathfrak{h}}^{+}\subsetneq\operatorname{\mathfrak{h}}^{+}$ for each $\operatorname{\mathfrak{h}}\in\mathcal{H}$. Suppose that there exists a hyperplane $\operatorname{\mathfrak{k}}$ of $Y$ which is not skewered by $g$. By [CS11, Proposition 3.2], every hyperplane of $Y$ is skewered by an element of $\Gamma$ so there exists $k\in\Gamma$ be such that $k\operatorname{\mathfrak{k}}^{+}\subsetneq\operatorname{\mathfrak{k}}^{+}$. By Lemma 2.3, there exists $n_{1}\in\mathbb{N}$ such that $g^{n_{1}}\operatorname{\mathfrak{k}}^{+}=\operatorname{\mathfrak{k}}^{+}$. Moreover, there exists $m\in\mathbb{N}$ such that, for each $\operatorname{\mathfrak{h}}\in\mathcal{H}$, either $k^{m}\operatorname{\mathfrak{h}}^{+}=\operatorname{\mathfrak{h}}^{+}$, $k^{m}\operatorname{\mathfrak{h}}^{+}\subsetneq\operatorname{\mathfrak{h}}^{+}$ or $k^{m}\operatorname{\mathfrak{h}}^{-}\subsetneq\operatorname{\mathfrak{h}}^{-}$. In the first two cases, $g^{n_{1}n}k^{m}\operatorname{\mathfrak{h}}^{+}\subseteq g^{n_{1}n}\operatorname{\mathfrak{h}}^{+}\subsetneq\operatorname{\mathfrak{h}}^{+}$ for all $n\in\mathbb{N}$. In the latter case, note that $d_{Haus}(\operatorname{\mathfrak{h}},k^{m}\operatorname{\mathfrak{h}})<\infty$ by Lemma 5.1 and, for any $n\in\mathbb{N}$ such that $d(g^{nn_{1}}\operatorname{\mathfrak{h}},\operatorname{\mathfrak{h}})>d_{Haus}(\operatorname{\mathfrak{h}},k^{m}\operatorname{\mathfrak{h}})$, we have that $g^{n_{1}n}k^{m}\operatorname{\mathfrak{h}}^{+}\subsetneq\operatorname{\mathfrak{h}}^{+}$. Since there are finitely many hyperplanes in $\mathcal{H}$, we can choose $n_{2}\in\mathbb{N}$ sufficiently large so that $g^{n_{1}n_{2}}k^{m}\operatorname{\mathfrak{h}}^{+}\subsetneq\operatorname{\mathfrak{h}}^{+}$ for all $\operatorname{\mathfrak{h}}\in\mathcal{H}$. We also have that $g^{n_{1}n_{2}}k^{m}\operatorname{\mathfrak{k}}^{+}\subsetneq g^{n_{1}n_{2}}\operatorname{\mathfrak{k}}^{+}=\operatorname{\mathfrak{k}}^{+}$ so the element $g^{\prime}\coloneqq g^{n_{1}n_{2}}k^{m}$ skewers every hyperplane in $\mathcal{H}^{\prime}\coloneqq\mathcal{H}\cup\{\operatorname{\mathfrak{k}}\}$. Moreover, for each $\operatorname{\mathfrak{h}}\in\mathcal{H}$, $\operatorname{\mathfrak{k}}$ cannot separate hyperplanes in $\operatorname{\langle}g\operatorname{\rangle}\cdot\operatorname{\mathfrak{h}}$ without being skewered by $g$. By the lack of facing triples, this implies that $\operatorname{\mathfrak{k}}$ must intersect $g^{n}\operatorname{\mathfrak{h}}$ for some $n$. It then follows from Lemma 5.1 that $\operatorname{\mathfrak{k}}$ intersects $g^{n}\operatorname{\mathfrak{h}}$ for all $n$. In particular, $\operatorname{\mathfrak{k}}$ intersects every hyperplane in $\mathcal{H}$, so $\mathcal{H}^{\prime}$ is pairwise intersecting and $|\mathcal{H}^{\prime}|\leq\dim(Y)$. If there exists a hyperplane $\operatorname{\mathfrak{k}}^{\prime}$ which is not skewered by $g^{\prime}$, then the above argument applied to $\mathcal{H}^{\prime},g^{\prime}$ and $\operatorname{\mathfrak{k}}^{\prime}$ implies that ′′$\mathcal{H}^{\prime\prime}\coloneqq\mathcal{H}^{\prime}\cup\{\operatorname{\mathfrak{k}}^{\prime}\}$ is pairwise intersecting and there exists $g^{\prime\prime}\in\Gamma$ which skewers every hyperplane in $\mathcal{H}^{\prime\prime}$. We can only iterate this process at most $\dim(Y)-1$ times until we obtain the desired element $\gamma$. ∎
@@
is well defined and, by [Cor16, Proposition 4.H.1], it is a homomorphism. Since $tr_{M^{+}}(\gamma)<0$, this homomorphism is non-trivial.
Clearly $K\leq\ker(tr_{M^{+}})$. Let us show that $K$ has finite index in $\ker(tr_{M^{+}})$. Given a hyperplane $\operatorname{\mathfrak{k}}$ of $Y$, define $L_{\operatorname{\mathfrak{k}}}\in\mathbb{Z}$ as follows. If $\operatorname{\mathfrak{k}}$ does not separate $\gamma^{n_{1}}\operatorname{\mathfrak{h}}$ from $\gamma^{n_{2}}\operatorname{\mathfrak{h}}$ for any $n_{1},n_{2}\in\mathbb{Z}$ then $L_{\operatorname{\mathfrak{k}}}\coloneqq 0$. Otherwise
@@
It remains to show that $\Gamma^{\prime}$ has finite index in $\Gamma$. Let $R\coloneqq d_{Haus}(\operatorname{\mathfrak{h}},\gamma\operatorname{\mathfrak{h}})<\infty$, let $\ell$ be a combinatorial geodesic axis for $\gamma$ and let $\sigma\subseteq\ell$ be a segment of length $R$. Let $1\leq R^{\prime}\leq R$ be such that exactly $R^{\prime}$ of the edges of $\sigma$ are dual to hyperplanes in the $\Gamma$-orbit of $\operatorname{\mathfrak{h}}$ and let $g_{1},\dots,g_{R^{\prime}}\in\Gamma$ be such that each hyperplane in $\Gamma\cdot\operatorname{\mathfrak{h}}$ which intersects $\sigma$ is of the form $g_{i}\operatorname{\mathfrak{h}}$ for some $i$. Let $g\in\Gamma$. Since $\gamma$ skewers every hyperplane in $Y$, the combinatorial axis $\ell$ intersects $g\gamma^{n}\operatorname{\mathfrak{h}}$ for all $n\in\mathbb{Z}$. Therefore the complement $\ell-\cup\{\ell\cap g\gamma^{n}\operatorname{\mathfrak{h}}:n\in\mathbb{Z}\}$ is a disjoint union of line segments of length at most $R$ which implies that, for some $n\in\mathbb{Z}$, the hyperplane $g\gamma^{n}\operatorname{\mathfrak{h}}$ intersects $\sigma$. Therefore there exists $i\in\{1,\dots,R^{\prime}\}$ such that $g\gamma^{n}\operatorname{\mathfrak{h}}=g_{i}\operatorname{\mathfrak{h}}$ which implies that $g_{i}^{-1}g\gamma^{n}\in K$ and $g_{i}^{-1}g\in K\gamma^{-n}\subseteq\Gamma^{\prime}$. Hence $\Gamma^{\prime}$ has index at most $R^{\prime}$ in $\Gamma$. ∎
@@
If we assume, as in the proof of Proposition 4.2, that $\Gamma$ and $H\coloneqq\operatorname{Stab}_{\Gamma}(\operatorname{\mathfrak{h}})$ are finitely generated and $H\operatorname{\setminus}\Gamma$ is narrow, then there is a much shorter proof:
For each $n\in\mathbb{N}$ let $\mathcal{H}_{n}$ denote the set of hyperplanes whose distance from $\operatorname{\mathfrak{h}}$ is exactly $n$. Then the hyperplanes of $\mathcal{H}_{n}$ contained in $\operatorname{\mathfrak{h}}^{-}$ are pairwise intersecting, as are those contained in $\operatorname{\mathfrak{h}}^{+}$. Therefore $|\mathcal{H}_{n}|\leq 2\dim(Y)$ so the stabiliser of any hyperplane in $\mathcal{H}_{n}$ with respect to the action of $H$ on $\mathcal{H}_{n}$ has index at most $(2\dim(Y))!$ in $H$. Since $H$ is finitely generated, there are finitely many such subgroups. By [CS11, Proposittion 3.2], there exists $g\in\Gamma$ such that $g\operatorname{\mathfrak{h}}^{+}\subsetneq\operatorname{\mathfrak{h}}^{+}$. The subgroup $K_{1}\coloneqq\cap_{n\in\mathbb{Z}}\operatorname{Stab}_{H}(g^{n}\operatorname{\mathfrak{h}})\leq H$ is a finite intersection of subgroups of index at most $(2\dim(Y))!$ so it is a finite index subgroup of $H$. Let $K\leq K_{1}$ be a finite index normal subgroup of $H$. Then for each $n\in\mathbb{Z}$, the index of $K^{g^{n}}$ in $K_{1}^{g^{n}}=K_{1}$ is equal to the index of $K$ in $K_{1}$ so the orbit $\{K^{g^{n}}:n\in\mathbb{Z}\}$ is finite. Therefore there exists $m\in\mathbb{N}$ such that $\operatorname{\langle}g^{m}\operatorname{\rangle}\leq N(K)$. For all $n\in\mathbb{N}$ and $k\in K$, $d(\operatorname{\mathfrak{h}},kg^{n}\operatorname{\mathfrak{h}})=d(\operatorname{\mathfrak{h}},g^{n}\operatorname{\mathfrak{h}})>d(\operatorname{\mathfrak{h}},g^{n-1}\operatorname{\mathfrak{h}})$ so each element of $\operatorname{\langle}g^{m}\operatorname{\rangle}$ is in a different coset of $K$ and $K$ has infinite index in its normaliser. Since $H\leq N(K)$ and $H\operatorname{\setminus}\Gamma$ is narrow, Proposition 3.6 then implies that $[\Gamma:N(K)]<\infty$.
The quotient group $N(K)/K$, equipped with any proper word metric, is quasi-isometric to $K\operatorname{\setminus}\Gamma$, which is a finite index cover of $H\operatorname{\setminus}\Gamma$. It is clear from the definition of $K$ that $K\leq\operatorname{Stab}_{\Gamma}(\operatorname{\mathfrak{h}}^{+})$ so $K\operatorname{\setminus}\Gamma$ is also a finite index cover of $\operatorname{Stab}_{\Gamma}(\operatorname{\mathfrak{h}}^{+})\operatorname{\setminus}\Gamma$. By [Sag95, §3.3], $\operatorname{Stab}_{\Gamma}(\operatorname{\mathfrak{h}}^{+})\operatorname{\setminus}\Gamma$ has at least two ends and, since it is narrow, $H\operatorname{\setminus}\Gamma$ has finitely many ends. It follows that $N(K)/K$ has $2\leq e<\infty$ ends but, since $N(K)/K$ is a group, $e$ must equal $2$ and $N(K)/K$ must be virtually cyclic [Hop44].